The U.S. Bureau of Labor Statistics released hourly wage figures for various countries for workers in the manufacturing sector. The hourly wage was $30.67 for Switzerland, $20.20 for Japan, and $23.82 for the U.S. Assume that in all three countries, the standard deviation of hourly labor rates is $4.00.

Appendix A Statistical Tables



a. Suppose 40 manufacturing workers are selected randomly from across Switzerland and asked what their hourly wage is. What is the probability that the sample average will be between $30.00 and $31.00?
Answer #1

Here, μ = 30.67, σ = 4/sqrt(40) = 0.6325, x1 = 30 and x2 = 31.

z = (x - μ)/σ
z1 = (30 - 30.67)/0.6325 = -1.06
z2 = (31 - 30.67)/0.6325 = 0.52

Therefore, we get
P(30 <= X <= 31) = P((31 - 30.67)/0.6325) <= z <= (31 - 30.67)/0.6325)
= P(-1.06 <= z <= 0.52) = P(z <= 0.52) - P(z <= -1.06)
= 0.6985 - 0.1446
= 0.5539